# Reasoning Models Are More Easily Gaslighted Than You Think

### v1

@@ Abstract.
Recent advances in reasoning-centric models promise improved robustness through mechanisms such as chain-of-thought prompting and test-time scaling. However, their ability to withstand misleading user input remains underexplored. In this paper, we conduct a systematic evaluation of three state-of-the-art reasoning models, i.e., OpenAI’s o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash, across three multimodal benchmarks: MMMU, MathVista, and CharXiv. Our evaluation reveals significant accuracy drops (25–29% on average) following gaslighting negation prompts, indicating that even top-tier reasoning models struggle to preserve correct answers under manipulative user feedback. Built upon the insights of the evaluation and to further probe this vulnerability, we introduce GaslightingBench-R, a new diagnostic benchmark specifically designed to evaluate reasoning models’ susceptibility to defend their belief under gaslighting negation prompt. Constructed by filtering and curating 1,025 challenging samples from the existing benchmarks, GaslightingBench-R induces even more dramatic failures, with accuracy drops exceeding 53% on average. Our findings reveal fundamental limitations in the robustness of reasoning models, highlighting the gap between step-by-step reasoning and belief persistence. Additional details are available on our project page: https://binzhubz.github.io/GaslightingBench-R/.
Reasoning models, Gaslighting negation, Multimodal reasoning
conference: ; ; copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
@@
Figure 1. (a) A qualitative example illustrating gaslighting-induced belief reversal. The model initially provides a correct, detailed explanation identifying four people wearing hats. After receiving a simple negation prompt, it revises the answer to five, hallucinating a new rationale. (b) Average accuracy of three reasoning models, OpenAI o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash on MMMU, MathVista, and CharXiv benchmarks, before and after gaslighting negation prompts. More detailed results are provided in Table 1.
@@
Recent advancements in Test-time scaling (Snell et al., 2025; Cobbe et al., 2021; Muennighoff et al., 2025) and chain-of-thought (Wei et al., 2022; Wang et al., 2023b) have significantly enhanced the reasoning capabilities of both Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs). Contemporary reasoning models, such as OpenAI’s o-series (Jaech et al., 2024) (e.g., o1, o3, o4-mini), Google’s Gemini-2.5-Flash (Team et al., 2023), Anthropic’s Claude-3.7-Sonnet (Anthropic, 2025), and DeepSeek-R1 (Guo et al., 2025), demonstrate impressive performance across a range of complex reasoning benchmarks, including mathematics, code generation, multimodal inference, and agentic tool use. These models are explicitly designed to “think deeper” through multi-step reasoning and typically require increased computational overhead at inference time. This distinguishes them from earlier non-reasoning models such as GPT-4o (Hurst et al., 2024), Qwen-VL (Bai et al., 2025), and LLaVA (Liu et al., 2024b), which prioritize speed and general-purpose instruction following, but often lack emergent “Aha moments” (Guo et al., 2025) characteristic of advanced reasoning models.
@@
Building on these advancements, many would argue that explicit reasoning mechanisms should enhance model robustness. By exposing intermediate steps, such as chain-of-thought traces, reasoning models are expected to self-inspect and, in principle, self-correct their reasoning before committing to a final answer (Miao et al., 2024). In theory, such transparency should offer a safeguard against prompt-based manipulation. Nevertheless, recent studies suggest that both LLMs and MLLMs remain surprisingly vulnerable to misleading user input. For instance, LLMs often exhibit sycophantic tendencies, ie., agreeing with user assertions even at the cost of factual accuracy due to biases introduced by human preference data (Sharma et al., 2024; Wang et al., [n. d.]). Similarly, as shown in Figure 1 (a), MLLMs have been shown to revise correct responses when confronted with gaslighting-style false claims (Zhu et al., 2025; Jiao et al., 2025). Despite these insights, the robustness of reasoning-centric models remains insufficiently understood. Given their explicit design for multi-step inference and “Aha moment” generation, one might anticipate greater resistance to manipulative prompts. Yet whether these structural features translate into genuine robustness remains an open and pressing question.
@@
To address this gap, we first conduct a comprehensive evaluation of state-of-the-art reasoning models with multimodal capabilities, including OpenAI’s o4-mini, Gemini-2.5-Flash (thinking mode), and Claude-3.7-Sonnet. Following OpenAI’s evaluation protocol, we assess multimodal reasoning performance using three established benchmarks: MMMU (Yue et al., 2024), MathVista (Lu et al., 2024) and CharXiv (Wang et al., 2024). For each dataset, our evaluation pipeline mirrors real-world user interactions by following (Zhu et al., 2025): both the question and its associated image are provided to the model. If a model’s initial answer is correct, we challenge it with a plausible yet incorrect user gaslighting negation to test whether the model maintains its original correct reasoning or revises its response. Furthermore, we introduce GaslightingBench-R (where “R” stands for Reasoning), a new benchmark specifically designed to evaluate the vulnerability of reasoning models to adversarial gaslighting prompts. GaslightingBench-R is constructed by curating and filtering representative samples from the three evaluated datasets, with the goals of maximizing domain diversity and preserving tasks that require genuine reasoning. The final benchmark covers 21 categories and includes 1,025 samples, encompassing a broad range of reasoning scenarios such as math, visual logic and chart.
@@
  • We conduct the first systematic study of gaslighting vulnerability in reasoning models, evaluating models including OpenAI-o4-mini, Gemini-2.5-Flash and Claude-3.7-Sonnet on three high-quality multimodal benchmarks: MMMU, MathVista, and ChartXiv.
  • We introduce GaslightingBench-R, a new benchmark specifically designed to assess the robustness of reasoning models under gaslighting negation prompt. The dataset is constructed from diverse and representative samples curated to test reasoning integrity across multiple domains.
@@
2. Related Work
2.1. Reasoning Models
@@
Large language models (LLMs) (Brown et al., 2020; Bai et al., 2023; Liu et al., 2024a; Grattafiori et al., 2024) and Multimodal Large Language Models (MLLMs) (Bai et al., 2025; Liu et al., 2024b; Hurst et al., 2024; Team et al., 2023) have driven significant progress in both natural language and multimodal understanding and generation. Building on this foundation, a new class of reasoning-centric models has emerged, distinguished by their explicit focus on test-time scaling (Snell et al., 2025; Cobbe et al., 2021; Muennighoff et al., 2025) and chain-of-thought (CoT) (Wei et al., 2022; Wang et al., 2023b) to improve the complex reasoning capabilities. CoT prompting encourages models to articulate intermediate reasoning steps prior to delivering a final answer, thereby mimicking human deductive processes. This approach has shown substantial improvements on tasks involving arithmetic reasoning, symbolic logic, and commonsense inference (Wei et al., 2022). The CoT paradigm has evolved rapidly with more advanced variants such as self-consistency with majority voting (Wang et al., 2023b), Least-to-Most Prompting (Zhou et al., 2023) and ReAct (Yao et al., 2023). Complementing CoT, test-time scaling (Snell et al., 2025; Cobbe et al., 2021; Muennighoff et al., 2025) dynamically adjusts computational resources during inference to support deeper thinking, such as longer context windows, deeper decoding, or iterative refinement. This allows models to “think harder” when needed, improving performance on complex and ambiguous tasks. Recently, the success of DeepSeek-R1 (Guo et al., 2025) via reinforcement learning (Sutton et al., 1998) has sparked growing interest in adapting similar training paradigms to multimodal reasoning tasks (Liu et al., 2025; Zhou et al., 2025; Shen et al., 2025; Huang et al., 2025). This paper complements this growing line of research by shifting the focus from performance on isolated reasoning tasks to robustness under adversarial gaslighting negation prompts. Specifically, we assess whether reasoning models can maintain correct beliefs when confronted with incorrect gaslighting negation. It is worth noting that many open-source multimodal reasoning models are currently based on relatively small architectures (e.g., 7B (Huang et al., 2025; Liu et al., 2025) or 2B (Zhou et al., 2025)) and exhibit limited interactive capabilities. As such, they are not well-suited for adversarial dialogue evaluations like ours. Consequently, we focus our study on state-of-the-art proprietary reasoning models that demonstrate stronger interaction competence and multimodal understanding.
2.2. Negation Understanding
@@
Negation is a fundamental linguistic concept that reverses semantic polarity and and plays a critical role in daily communication. Despite their impressive performance on a wide range of NLP tasks, large pre-trained language models such as BERT, GPT-3, and InstructGPT have been shown to struggle with negation comprehension (Ettinger, 2020; Kassner and Schütze, 2020; Truong et al., 2023). More recently, studies have demonstrated that even advanced LLMs fail to defend correct beliefs when challenged with invalid opposing statement (Wang et al., [n. d.]; Zhao et al., 2025), highlighting a persistent vulnerability in adversarial deceptive contexts. In multimodal settings, negation understanding becomes more complex due to the need to ground language in visual perception. Existing works (Wang et al., 2023a; Yuksekgonul et al., 2023; Alhamoud et al., 2025) have shown that CLIP-like vision-language models perform poorly on negation-sensitive tasks like cross-modal retrieval and visual question answering. These limitations are often attributed to the models’ inability to encode and reason about negated visual concepts, though recent efforts suggest that performance can be improved with negation-oriented datasets and contrastive fine-tuning (Alhamoud et al., 2025). Our prior GaslightingBench (Zhu et al., 2025) demonstrates that even strong MLLMs are highly susceptible to negations and significantly overturn the correct answers under adversarial gaslighting negation prompts, which can be mitigated by attention reallocation (Jiao et al., 2025). Different from GaslightingBench (Zhu et al., 2025), this paper aims to specifically target the latest class of reasoning models that combine chain‐of‐thought with test‐time scaling. We comprehensively evaluate the impact of gaslighting negation on reasoning models and build a new benchmark GaslightingBench-R to assess reasoning stability in multimodal contexts.
3. GaslightingBench-R: Gaslighting Benchmark for Reasoning Models
3.1. Evaluating Reasoning Models with Gaslighting Negation Prompt
@@
Reasoning Models. We focus our evaluation on three state-of-the-art reasoning models that exhibit remarkable multimodal capabilities: OpenAI’s o4-mini (Jaech et al., 2024), Google’s Gemini-2.5-Flash (thinking mode) (Team et al., 2023), and Anthropic’s Claude-3.7-Sonnet (Anthropic, 2025). These models are explicitly optimized for advanced reasoning through mechanisms such as chain-of-thought prompting and increased test-time computation. They have achieved state-of-the-art performance across a range of benchmarks involving mathematics, code generation, and visual inference. Each model is accessed and evaluated via its publicly available API under standardized settings.
@@
Benchmarks. To evaluate the robustness of these reasoning models under gaslighting negation prompt, we use three high-quality multimodal reasoning datasets following OpenAI o-series (Jaech et al., 2024): MMMU (Yue et al., 2024), MathVista (Lu et al., 2024) and CharXiv (Wang et al., 2024). MMMU is a challenging benchmark composed of university-level exam question-image pairs spanning diverse disciplines such as biology, physics, and history, often requiring domain knowledge and visual comprehension. MathVista focuses on complex mathematical reasoning in multimodal settings, combining textual descriptions with visual aids such as plots, geometric figures, and tables. CharXiv derives from arXiv papers, consists of real-world scientific charts and graphs paired with questions that require advanced reasoning and comprehension of quantitative visual data.
Figure 2. GaslightingBench-R spans a diverse set of categories drawn from three multimodal reasoning benchmarks (MMMU, MathVista, and CharXiv).
@@
Evaluation Pipeline. Our evaluation follows a two-stage framework inspired by (Zhu et al., 2025). For each sample, the model is presented with the question and corresponding image. If the model’s initial response matches the ground truth, it is deemed correct and proceeds to the gaslighting phase. In this second stage, we introduce a gaslighting prompt, a confidently phrased negation of the correct answer (e.g., “No, that’s incorrect. Please verify your answer.”) designed to simulate misleading user behavior in a realistic conversational setting. We then observe whether the model maintains its original response or revises it in light of the adversarial intervention. A successful gaslighting instance is recorded when a model changes a previously correct answer to an incorrect one. This evaluation framework allows us to assess a model’s consistency and robustness to manipulative dialogue. To ensure fair comparison across models, we adopt a standardized conversational prompting template throughout all experiments.
3.2. Reasoning Gaslighting Benchmark Construction

### v2

@@ Abstract
Recent advances in reasoning-centric models promise improved robustness through mechanisms such as chain-of-thought prompting and test-time scaling. However, their ability to withstand gaslighting negation attacks—adversarial prompts that confidently deny correct answers-remains underexplored. In this paper, we conduct a systematic evaluation of three state-of-the-art reasoning models, i.e., OpenAI’s o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash, across three multimodal benchmarks: MMMU, MathVista, and CharXiv. Our evaluation reveals significant accuracy drops (25–29% on average) following gaslighting negation attacks, indicating that even top-tier reasoning models struggle to preserve correct answers under manipulative user feedback. Built upon the insights of the evaluation and to further probe this vulnerability, we introduce GaslightingBench-R, a new diagnostic benchmark specifically designed to evaluate reasoning models’ susceptibility to defend their belief under gaslighting negation attacks. Constructed by filtering and curating 1,025 challenging samples from the existing benchmarks, GaslightingBench-R induces even more dramatic failures, with accuracy drops exceeding 53% on average. Our findings highlight a fundamental gap between step-by-step reasoning and resistance to adversarial manipulation, calling for new robustness strategies that safeguard reasoning models against gaslighting negation attacks. Additional details are available on our project page: https://binzhubz.github.io/GaslightingBench-R/.
11footnotetext: $\dagger$ Jingjing Chen is the corresponding author.22footnotetext: * Equal contribution.
1 Introduction
@@
Figure 1: (a) A qualitative example illustrating gaslighting negation attack. The model initially provides a correct, detailed explanation identifying four people wearing hats. After receiving a simple negation prompt, it revises the answer to five, hallucinating a new rationale. (b) Average accuracy of three reasoning models, OpenAI o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash on MMMU, MathVista, and CharXiv benchmarks, before and after gaslighting negation attacks. More detailed results are provided in Table 1.
@@
Recent advancements in Test-time scaling [snell2025scaling, cobbe2021training, muennighoff2025s1] and chain-of-thought [wei2022chain, wangself] have significantly enhanced the reasoning capabilities of both Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs). Contemporary reasoning models, such as OpenAI’s o-series [jaech2024openai] (e.g., o1, o3, o4-mini), Google’s Gemini-2.5-Flash [team2023gemini], Anthropic’s Claude-3.7-Sonnet [anthropic2025claude37], and DeepSeek-R1 [guo2025deepseek], demonstrate impressive performance across a range of complex reasoning benchmarks, including mathematics, code generation, multimodal inference, and agentic tool use. These models are explicitly designed to “think deeper” through multi-step reasoning and typically require increased computational overhead at inference time. This distinguishes them from earlier non-reasoning models such as GPT-4o [hurst2024gpt], Qwen-VL [bai2025qwen2], and LLaVA [liu2024improved], which prioritize speed and general-purpose instruction following, but often lack emergent “Aha moments” [guo2025deepseek] characteristic of advanced reasoning models.
@@
Building on these advancements, many would argue that explicit reasoning mechanisms should enhance model robustness. By exposing intermediate steps, such as chain-of-thought traces, reasoning models are expected to self-inspect and, in principle, self-correct their reasoning before committing to a final answer [miaoselfcheck]. In theory, such transparency should offer a safeguard against prompt-based manipulation. Nevertheless, recent studies suggest that both LLMs and MLLMs remain surprisingly vulnerable to misleading user input. For instance, LLMs often exhibit sycophantic tendencies, ie., agreeing with user assertions even at the cost of factual accuracy due to biases introduced by human preference data [sharmatowards, wang2023can]. Similarly, as shown in Figure 1 (a), MLLMs have been shown to revise correct responses when confronted with gaslighting-style false claims [zhu2025calling, jiao2025don, SpeechGaslighting]. Despite these insights, the robustness of reasoning-centric models remains insufficiently understood. Given their explicit design for multi-step inference and “Aha moment” generation, one might anticipate greater resistance to manipulative prompts. Yet whether these structural features translate into genuine robustness remains an open and pressing question.
@@
To address this gap, we first conduct a comprehensive evaluation of state-of-the-art reasoning models with multimodal capabilities, including OpenAI’s o4-mini, Gemini-2.5-Flash (thinking mode), and Claude-3.7-Sonnet. Following OpenAI’s evaluation protocol, we assess multimodal reasoning performance using three established benchmarks: MMMU [mmmu], MathVista [mathvista] and CharXiv [wang2024charxiv]. For each dataset, our evaluation pipeline mirrors real-world user interactions by following [zhu2025calling]: both the question and its associated image are provided to the model. If a model’s initial answer is correct, we challenge it with a plausible yet incorrect user gaslighting negation to test whether the model maintains its original correct reasoning or revises its response. Furthermore, we introduce GaslightingBench-R (where “R” stands for Reasoning), a new benchmark specifically designed to evaluate the vulnerability of reasoning models to gaslighting negation attacks. GaslightingBench-R is constructed by curating and filtering representative samples from the three evaluated datasets, with the goals of maximizing domain diversity and preserving tasks that require genuine reasoning. The final benchmark covers 21 categories and includes 1,025 samples, encompassing a broad range of reasoning scenarios such as math, visual logic and chart.
@@
  • We conduct the first systematic study of gaslighting negation attack in reasoning models, evaluating models including OpenAI-o4-mini, Gemini-2.5-Flash and Claude-3.7-Sonnet on three high-quality multimodal benchmarks: MMMU, MathVista, and ChartXiv.
  • We introduce GaslightingBench-R, a new benchmark specifically designed to assess the robustness of reasoning models under gaslighting negation attack. The dataset is constructed from diverse and representative samples curated to test reasoning integrity across multiple domains.
@@
2 Related Work
2.1 Reasoning Models
@@
Large language models (LLMs) [brown2020language, bai2023qwen, liu2024deepseek, grattafiori2024llama] and Multimodal Large Language Models (MLLMs) [bai2025qwen2, liu2024improved, hurst2024gpt, team2023gemini, jiao2025holistic] have driven significant progress in both natural language and multimodal understanding and generation. Building on this foundation, a new class of reasoning-centric models has emerged, distinguished by their explicit focus on test-time scaling [snell2025scaling, cobbe2021training, muennighoff2025s1] and chain-of-thought (CoT) [wei2022chain, wangself] to improve the complex reasoning capabilities. CoT prompting encourages models to articulate intermediate reasoning steps prior to delivering a final answer, thereby mimicking human deductive processes. This approach has shown substantial improvements on tasks involving arithmetic reasoning, symbolic logic, and commonsense inference [wei2022chain]. The CoT paradigm has evolved rapidly with more advanced variants such as self-consistency with majority voting [wangself], Least-to-Most Prompting [zhouleast] and ReAct [yaoreact]. Complementing CoT, test-time scaling [snell2025scaling, cobbe2021training, muennighoff2025s1] dynamically adjusts computational resources during inference to support deeper thinking, such as longer context windows, deeper decoding, or iterative refinement. This allows models to “think harder” when needed, improving performance on complex and ambiguous tasks. Recently, the success of DeepSeek-R1 [guo2025deepseek] via reinforcement learning [sutton1998reinforcement] has sparked growing interest in adapting similar training paradigms to multimodal reasoning tasks [liu2025visual, zhou2025r1, shen2025vlm, huang2025vision]. This paper complements this growing line of research by shifting the focus from performance on isolated reasoning tasks to robustness under adversarial gaslighting negation attacks. Specifically, we assess whether reasoning models can maintain correct beliefs when confronted with incorrect gaslighting negation. It is worth noting that many open-source multimodal reasoning models are currently based on relatively small architectures (e.g., 7B [huang2025vision, liu2025visual] or 2B [zhou2025r1]) and exhibit limited interactive capabilities. As such, they are not well-suited for adversarial dialogue evaluations like ours. Consequently, we focus our study on state-of-the-art proprietary reasoning models that demonstrate stronger interaction competence and multimodal understanding.
2.2 Negation Understanding
@@
Negation is a fundamental linguistic concept that reverses semantic polarity and and plays a critical role in daily communication. Despite their impressive performance on a wide range of NLP tasks, large pre-trained language models such as BERT, GPT-3, and InstructGPT have been shown to struggle with negation comprehension [ettinger2020bert, kassner2020negated, truong2023language]. More recently, studies have demonstrated that even advanced LLMs fail to defend correct beliefs when challenged with invalid opposing statement [wang2023can, zhao2025aligning], highlighting a persistent vulnerability in adversarial deceptive contexts. In multimodal settings, negation understanding becomes more complex due to the need to ground language in visual perception. Existing works [wang2023clipn, yuksekgonuland, alhamoud2025vision] have shown that CLIP-like vision-language models perform poorly on negation-sensitive tasks like cross-modal retrieval and visual question answering. These limitations are often attributed to the models’ inability to encode and reason about negated visual concepts, though recent efforts suggest that performance can be improved with negation-oriented datasets and contrastive fine-tuning [alhamoud2025vision]. The prior GaslightingBench [zhu2025calling] demonstrates that even strong MLLMs are highly susceptible to negations and significantly overturn the correct answers under adversarial gaslighting negation attacks, which can be mitigated by attention reallocation [jiao2025don]. Different from GaslightingBench [zhu2025calling], this paper aims to specifically target the latest class of reasoning models that combine chain‐of‐thought with test‐time scaling. We comprehensively evaluate the impact of gaslighting negation attack on reasoning models and build a new benchmark GaslightingBench-R to assess reasoning stability in multimodal contexts.
3 GaslightingBench-R: Gaslighting Benchmark for Reasoning Models
3.1 Evaluating Reasoning Models with Gaslighting Negation Attack
@@
Reasoning Models. We focus our evaluation on three state-of-the-art reasoning models that exhibit remarkable multimodal capabilities: OpenAI’s o4-mini [jaech2024openai], Google’s Gemini-2.5-Flash (thinking mode) [team2023gemini], and Anthropic’s Claude-3.7-Sonnet [anthropic2025claude37]. These models are explicitly optimized for advanced reasoning through mechanisms such as chain-of-thought prompting and increased test-time computation. They have achieved state-of-the-art performance across a range of benchmarks involving mathematics, code generation, and visual inference. Each model is accessed and evaluated via its publicly available API under standardized settings. While open-source multimodal reasoning models are rapidly emerging, most are still limited to relatively small-scale architectures (e.g., 2B–7B parameters [huang2025vision, liu2025visual, zhou2025r1]) and exhibit restricted interactive dialogue capabilities. These constraints make them less suitable for stress-testing adversarial scenarios that require multi-turn engagement and belief persistence. We therefore center our study on advanced proprietary reasoning models that offer stronger interactive competence and richer multimodal understanding, allowing for a more realistic and rigorous assessment of robustness under gaslighting attacks. It is also worth noting that models such as DeepSeek-R1, DeepSeek-V3.1, gpt-oss, and QwQ-32B are restricted to text-only reasoning, and therefore fall outside the scope of our multimodal evaluation.
@@
Benchmarks. To evaluate the robustness of these reasoning models under gaslighting negation attacks, we then use three high-quality multimodal reasoning datasets following OpenAI o-series [jaech2024openai]: MMMU [mmmu], MathVista [mathvista] and CharXiv [wang2024charxiv]. MMMU is a challenging benchmark composed of university-level exam question-image pairs spanning diverse disciplines such as biology, physics, and history, often requiring domain knowledge and visual comprehension. MathVista focuses on complex mathematical reasoning in multimodal settings, combining textual descriptions with visual aids such as plots, geometric figures, and tables. CharXiv derives from arXiv papers, consists of real-world scientific charts and graphs paired with questions that require advanced reasoning and comprehension of quantitative visual data.
@@
Evaluation Pipeline. Our evaluation follows a two-stage framework inspired by [zhu2025calling]. For each sample, the model is presented with the question and corresponding image. If the model’s initial response matches the ground truth, it is deemed correct and proceeds to the gaslighting phase. In this second stage, we introduce a gaslighting prompt, a confidently phrased negation of the correct answer (e.g., “No, that’s incorrect. Please verify your answer.”) designed to simulate misleading user behavior in a realistic conversational setting. We then observe whether the model maintains its original response or revises it in light of the adversarial intervention. A successful gaslighting instance is recorded when a model changes a previously correct answer to an incorrect one. This evaluation framework allows us to assess a model’s consistency and robustness to manipulative dialogue. To ensure fair comparison across models, we adopt a standardized conversational prompting template throughout all experiments.
3.2 Reasoning Gaslighting Benchmark Construction
Figure 2: GaslightingBench-R spans a diverse set of categories drawn from three multimodal reasoning benchmarks (MMMU, MathVista, and CharXiv).
Figure 3: Examples from different categories in the GaslightingBench-R.